OpenStudy (anonymous):

what is the y intercept of h(x) = 2 cos(2x − π) + 4

OpenStudy (anonymous):

y-intercepts are coordinate pairs of the form \((0,y)\). So, plugging in \(x=0\) will give you the y-intercept: \[h(0)=2\cos(2(0)-\pi)+4=\cdots\]